OpenStudy (anonymous):

tan^2(x)=f(g9X)) what is f(x)= g(x), Please include process for defining I can find reference for this?

OpenStudy (anonymous):

Sorry shoudl be tan^2(x)=f(g(x))

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

It may be easier to see the composition if you write \(\tan^2x=\left(\tan x\right)^2\). Here, the inside function \((g(x))\) is \(\tan x\), and the outside function \((f(x))\) is \(x^2\). So, \(f(g(x))=f(\tan x)=\tan^2x\), like you wanted. Are you supposed to find \(f(x)+g(x)\) ?

OpenStudy (anonymous):

I am to define f(x) and g(x) I will input and verify

OpenStudy (anonymous):

No that did not work, I believe I am required to rewrite the original function Tan^2(x)=f(g(x)) in terms of f(x)= ? and g(x)= ? not sure just yet